### v6

@@ 5.3. Polynomial Polynomials
For our third case study, we provide sparse and dense representations of univariate polynomials with natural number coefficients, both of which are represented using setoids. The former representation uses the simplest possible data type to represent a polynomial, while the latter representation will mimic the way polynomials are commonly written, which will utilize far less storage for polynomials of high degree with many terms having a coefficient of 0. We implement and repair addition of polynomials and evaluation of polynomials on a natural number, as well as proofs that addition is commutative and that evaluation respects addition. Our proofs can be found in polynomial.v.
@@ 6. Correctness
Up to now, we have been imprecise with what it means for repair to be conducted correctly. The original Pumpkin Pi paper (Ringer et al., 2021) outlines a definition for correctness of repair using univalent type theory. We review that definition here, giving the needed background in univalent type theory. Then, we implement this definition in Cubical Agda, a proof assistant which has the necessary univalent type theory. This allows us to, for the first time, construct proofs that functions and theorems were correctly repaired.
To implement correctness of repair, we need a notion of heterogeneous equality. The Pumpkin Pi paper uses dependent path equality in a univalent type system for this purpose. Univalence states that equivalence is equivalent to equality (Cohen et al., 2018; Angiuli et al., 2017): any equivalence of types corresponds to a unique proof of equality between those two types. We term these equalities path equalities. Given a path equality p : A $\equiv$ B between two types and given elements a : A, b : B of those types, we can construct the type of dependent path equalities between those elements, notated PathP p a b. This type is inhabited if and only if the equivalence of types corresponding to p by univalence maps a to b.
@@
Thus, to use this approach to construct proofs of correct repair, we need to work in a type system that supports this notion of dependent path equality. Coq does not support univalence without adding axioms to its type system. Instead, we implement this approach using Cubical Agda, which has univalence as a theorem. As a technical detail, we restrict ourselves to working with types which are h-sets: that is, types where uniqueness of identity proofs holds. As a bonus, Cubical Agda allows a direct implementation of quotient types via the following higher inductive type:
@@
The first constructor is the constructor for an element of a quotient type as described in Section 2, and the second encodes the equality property for quotient types. The third constructor enforces that quotient types are h-sets. We will use this type instead of setoids when specifying correctness of repair in Cubical Agda. As a result, the additional rules to repair equivalence relations from Figure 5 in Section 3 are not needed when conducting repair in Cubical Agda. The transformation from Figure 4 in Section 3 works, assuming the same annotations as in Coq, noting that we cannot reuse any of Pumpkin Pi’s automation in Cubical Agda. There is only one additional restriction: for a quotient type Q, every motive P : Q $\to$ Set comes with the requirement that ((x : Q) $\to$ isSet (P x)), to ensure that we actually do stay in the h-set fragment of Cubical Agda. To demonstrate that our repair methodology still works under this paradigm, we manually followed the transformation to repair the functions and proofs from the first two case studies in Section 5, which can be found in the supplementary materials in files grothendieck_int_equiv.agda and equivalence_queue.agda respectively.
@@
$\forall$ (PA : $\mathbb{N}$to Type) (PB : Int / rInt $\to$ Type) (PBSet : $\forall$ b $\to$ isSet (PB b)) $\to$
@@
When constructing proofs of correct repair for theorems, because of difficulties when composing different PathPs, our rules for proving correct repair do not always compose directly in their current formulation. As a result, if Cubical Agda had the facilities for building automation, we would not currently have a complete procedure for constructing proofs of correct repair fully automatically. However, we are able to prove these theorems directly in such cases, demonstrating that our repair procedure produces the desired output. All of the rules we prove, as well as the proofs of correct repair in our example, can be found in equivalence_int_abs.agda in the supplementary material.
@@ 7. Related Work
Proof Repair. This work extends the Pumpkin P(Ringer et al., 2021) proof repair transformation and Coq plugin to support quotient type equivalences, a class of changes previously not supported. Pumpkin Pi has some more mature automation for other classes of changes, like automatic search for configurations, that we do not yet extend to work for quotient type equivalences. Proof repair was first introduced in parallel by Ringer et al. (2018) and Robert (2018), with strong influence from program repair (Monperrus, 2017). Sisyphus (Gopinathan et al., 2023) is a recent proof repair tool that, like our work, can handle changes in behavior (using a mix of dynamic and static techniques). However, Sisyphus repairs proofs of imperative OCaml programs verified in Coq using an embedded separation logic, whereas our work repairs proofs that are written in Coq directly.
Univalent Foundations. Parts of this project are grounded in Cubical Agda, and parts assume a univalent metatheory. Cubical Agda is an implementation of cubical type theory (Vezzosi et al., 2019). Cubical type theory (Cohen et al., 2018; Coquand et al., 2018; Angiuli et al., 2017) was developed to give a constructive account of the univalence axiom. When working in Cubical Agda, we are able to state and prove internal correctness of parts of our repair transformation and have a computational interpretation of functional extensionality. Cubical type theory itself is a derivative of Voevodsky’s homotopy type theory (Univalent Foundations Program, 2013), which presents the univalence axiom non-constructively. Homotopy type theory has additionally been implemented in Coq as the HoTT library (Bauer et al., 2017).
@@
Work has been done to implement transfer tools in Coq that approximate or externally implement automation corresponding to univalent transport. Tabareau et al. (2018) defines univalent parametricity, which allows transport of a restricted class of functions and theorems. Univalent parametricity implements an ad hoc form of transport that only sometimes requires functional extensionality, and in many cases is axiom-free. It also includes a form of type-directed search to transport terms by way of type classes, something that proof repair tools like Pumpkin Pi and our extension still lack. Subsequent work introduces a white-box transformation (Tabareau et al., 2021) similar to the repair transformation from Pumpkin Pi, which Ringer (2021) describes as developed in parallel with mutual influence. None of these support quotient type equivalences like our work does, though it is possible that by leaning further on functional extensionality, one could use these tools with quotient type equivalences.
@@
More recent work called, Trocq (Cohen et al., 2024) implements external transfer for Coq that directly supports relations more general than equivalences, like CoqEAL, but also supports proofs. Like Pumpkin Pi, Trocq goes out of its way to avoid depending on axioms like univalence and functional extensionality. Trocq’s motivation of supporting transfer of proofs across relations more general than equivalences is similar to the motivation of our extensions to Pumpkin Pi, with two differences: (1) our work supports a more limited class of relations that can be described as equivalences between quotient types, and (2) for that class of changes, by extending Pumpkin Pi’s proof term transformation, our work makes it possible to remove the old version of a type after applying repair. The major benefit of our tool relative to Trocq comes from (2)—while all proof repair tools implement a kind of transfer, not all transfer methods implement repair, and Trocq does not implement repair.
@@
Quotients and Equivalences. Our work uses quotient types to expand the scope of proof repair. Quotient types exist in other proofs assistants besides Cubical Agda, like Isabelle/HOL (Wenzel et al., 2004; Isabelle Development Team, 2024), as well as Lean (Avigad et al., 2017) by way of axioms. Bortin and Lüth (2010) use quotient types to construct theories in Isabelle, like multisets and finite sets as quotients of lists. Coq does not have quotient types, but it does have setoids (Sozeau, 2023), which do not explicitly form equivalence classes like quotients do. Setoid type theory uses a setoid model to justify the axioms needed to represent quotient types (Altenkirch et al., 2019). We draw on internal quotient types for our grounding in Cubical Agda, and we draw on setoids for our work in Coq.
@@ 8. Conclusions & Future Work
We extended Pumpkin Pi to support changes represented by quotient type equivalences, enabling repair in situations previously untenable. The key challenge we overcame was supporting quotient types in a proof repair algorithm built for a type theory that does not have quotient types to begin with. We addressed this by representing quotient types using setoids, extending the Pumpkin Pi algorithm and implementation to repair proofs about equivalence relations, and adding new automation to dispatch newly generated proof obligations. Our extension demonstrated success on three case studies not supported by the original Pumpkin Pi. We also constructed the first internal correctness proofs for repair. We wish to continue to improve our extension’s automation and usability, and we hope to look at other kinds of types and relations that can be expressed even when the type theory lacks them, as quotient types can be by way of setoids. We hope this will open the door to proof repair for more sophisticated classes of changes.

### v7

@@ 5.3. Polynomial Polynomials
For our third case study, we provide sparse and dense representations of univariate polynomials with natural number coefficients, both of which are represented using setoids. The former representation uses the simplest possible data type to represent a polynomial, while the latter representation will mimic the way polynomials are commonly written, which will utilize far less storage for polynomials of high degree with many terms having a coefficient of 0. We implement and repair addition of polynomials and evaluation of polynomials on a natural number, as well as proofs that addition is commutative and that evaluation respects addition. Our proofs can be found in the artifact   $1$.
@@ 6. Correctness
Up to now, we have been imprecise with what it means for repair to be conducted correctly. The original Pumpkin Pi paper (Ringer et al., 2021) outlines a definition for correctness of repair using univalent type theory. We review that definition here, giving the needed background in univalent type theory. Then, we implement this definition in a proof assistant which has the necessary univalent type theory. This allows us to, for the first time, construct proofs that functions and theorems were correctly repaired.
To state and prove correctness of repair, we need a notion of heterogeneous equality. The Pumpkin Pi paper uses dependent path equality in a univalent type system for this purpose. Univalence states that equivalence is equivalent to equality (Univalent Foundations Program, 2013; Cohen et al., 2018; Angiuli et al., 2017): any equivalence of types corresponds to a unique proof of equality between those two types. We term these equalities path equalities. Given a path equality p : A $\equiv$ B between types and given elements a : A, b : B of those types, we can construct the type of dependent path equalities between those elements, notated PathP p a b. This type is inhabited if and only if the equivalence of types corresponding to p by univalence maps a to b.
@@
Thus, to use this approach to construct proofs of correct repair, we need to work in a type system that supports this notion of dependent path equality. For this, we choose Cubical Agda. It is true that we could have just as well used Rocq’s HoTT library (Bauer et al., 2017), which treats univalence as an axiom. However, Rocq’s HoTT library also fundamentally changes how the Prop universe works. Switching to such a library for the sake of constructing correctness proofs would have imposed as much overhead for us as directly using a different proof assistant that has univalence, and may have been extra confusing to readers as it would have looked the same as Rocq while relying on fundamentally different dependencies that change the type theory. Therefore, we instead switch to Cubical Agda. We also like the bonus that terms in Cubical Agda compute, since Cubical Agda has univalence as a theorem rather than an axiom.
As a technical detail, we restrict ourselves to working with types which are h-sets: that is, types where uniqueness of identity proofs holds. As a bonus, Cubical Agda allows a direct implementation of quotient types via the following higher inductive type:
@@
The first constructor is the constructor for an element of a quotient type as described in Section 2, and the second encodes the equality property for quotient types. The third constructor enforces that quotient types are h-sets. We will use this type instead of setoids when specifying correctness of repair in Cubical Agda. As a result, the additional rules to repair equivalence relations from Figure 5 in Section 3 are not needed when conducting repair in Cubical Agda. The transformation from Figure 4 in Section 3 works, assuming the same annotations as in Rocq, noting that we cannot reuse any of Pumpkin Pi’s automation in Cubical Agda. There is only one additional restriction: for a quotient type Q, every motive P : Q $\to$ Set comes with the requirement that ((x : Q) $\to$ isSet (P x)), to ensure that we actually do stay in the h-set fragment of Cubical Agda. To demonstrate that our repair methodology still works under this paradigm, we manually followed the transformation to repair the functions and proofs from the first two case studies in Section 5, which can be found in the artifact   12   13.
@@
$\forall$ (PA : $\mathbb{N}$ $\to$ Type) (PB : Int / rInt $\to$ Type) (PBSet : $\forall$ b $\to$ isSet (PB b)) $\to$
@@
When constructing proofs of correct repair for theorems, because of difficulties when composing different PathPs, our rules for proving correct repair do not always compose directly in their current formulation. As a result, if Cubical Agda had the facilities for building automation, we would not currently have a complete procedure for constructing proofs of correct repair fully automatically. However, we are able to prove these theorems directly in such cases, demonstrating that our repair procedure produces the desired output. All of the rules we prove, as well as the proofs of correct repair in our example, can be found in the artifact   14.
@@ 7. Related Work
Proof Repair. This work extends the Pumpkin P(Ringer et al., 2021) proof repair transformation and Rocq plugin to support quotient type equivalences, a class of changes previously not supported. Pumpkin Pi has some more mature automation for other classes of changes, like automatic search for configurations, that we do not yet extend to work for quotient type equivalences. Proof repair was first introduced in parallel by Ringer et al. (2018) and Robert (2018), with strong influence from program repair (Monperrus, 2018). Sisyphus (Gopinathan et al., 2023) is a recent proof repair tool that, like our work, can handle changes in behavior (using a mix of dynamic and static techniques). However, Sisyphus repairs proofs of imperative OCaml programs verified in Rocq using an embedded separation logic, whereas our work repairs proofs that are written in Rocq directly.
Univalent Foundations. Parts of this project are grounded in Cubical Agda, and parts assume a univalent metatheory. Cubical Agda is an implementation of cubical type theory (Vezzosi et al., 2019). Cubical type theory (Cohen et al., 2018; Coquand et al., 2018; Angiuli et al., 2017) was developed to give a constructive account of the univalence axiom. When working in Cubical Agda, we are able to state and prove internal correctness of parts of our repair transformation and have a computational interpretation of functional extensionality. Cubical type theory itself is a derivative of Voevodsky’s homotopy type theory (Univalent Foundations Program, 2013), which presents the univalence axiom non-constructively. Homotopy type theory has additionally been implemented in Rocq as the HoTT library (Bauer et al., 2017). Univalence leads to a related idea, the Structure Identity Principle, which states that the type of equalities of structures is equivalent to the type of isomorphisms of those structures (Univalent Foundations Program, 2013; Coquand and Danielsson, 2013; Ahrens et al., 2022).
@@
Work has been done to implement transfer tools in Rocq that approximate or externally implement automation corresponding to univalent transport. Tabareau et al. (2018) defines univalent parametricity, which allows transport of a restricted class of functions and theorems. Univalent parametricity implements an ad hoc form of transport that only sometimes requires functional extensionality, and in many cases is axiom-free. It also includes a form of type-directed search to transport terms by way of type classes, something that proof repair tools like Pumpkin Pi and our extension still lack. Subsequent work introduces a white-box transformation (Tabareau et al., 2021) similar to the repair transformation from Pumpkin Pi, which Ringer (2021) describes as developed in parallel with mutual influence. None of these support quotient type equivalences like our work does, though it is possible that by leaning further on functional extensionality, one could use these tools with quotient type equivalences.
@@
More recent work called, Trocq (Cohen et al., 2024) implements external transfer for Rocq that directly supports relations more general than equivalences, like CoqEAL, but also supports proofs. Like Pumpkin Pi, Trocq goes out of its way to avoid depending on axioms like univalence and functional extensionality. Trocq’s motivation of supporting transfer of proofs across relations more general than equivalences is similar to the motivation of our extensions to Pumpkin Pi, with two differences: (1) our work supports a more limited class of relations that can be described as equivalences between quotient types, and (2) for that class of changes, by extending Pumpkin Pi’s proof term transformation, our work makes it possible to remove the old version of a type after applying repair. The major benefit of our tool relative to Trocq comes from (2)—while all proof repair tools implement a kind of transfer, not all transfer methods implement repair, and Trocq does not implement repair.
@@
Quotients and Equivalences. Our work uses quotient types (Hofmann, 1995) to expand the scope of proof repair. Maietti (1999) shows that, in intensional Martin-Löf type theory with uniqueness of identity proofs and at least two universes, the existence of effective quotients implies the law of the excluded middle for types in the first universe. Quotient types exist in other proofs assistants besides Cubical Agda, like Isabelle/HOL (Wenzel et al., 2004; Isabelle Development Team, 2024), as well as Lean (Avigad et al., 2017) by way of axioms. Bortin and Lüth (2010) use quotient types to construct theories in Isabelle, like multisets and finite sets as quotients of lists. Rocq does not have quotient types, but it does have setoids (Sozeau, 2023), which do not explicitly form equivalence classes like quotients do. Setoid type theory uses a setoid model to justify the axioms needed to represent quotient types (Altenkirch et al., 2019). XTT uses Bishop sets, analogous to setoids, to construct a type theory where all types have definitional uniqueness of identity proofs (Sterling et al., 2022). We draw on quotient types for our work in Cubical Agda, and we draw on setoids for our work in Rocq.
@@ 8. Conclusions & Future Work
We extended Pumpkin Pi to support changes represented by quotient type equivalences, enabling repair in situations previously untenable. The key challenge we overcame was supporting quotient types in a proof repair algorithm built for a type theory that does not have quotient types to begin with. We addressed this by representing quotient types using setoids, extending the Pumpkin Pi algorithm and implementation to repair proofs about equivalence relations, and adding new automation to dispatch newly generated proof obligations. Our extension demonstrated success on three case studies not supported by the original Pumpkin Pi. We also constructed the first internal correctness proofs for repair. We wish to continue to improve our extension’s automation and usability, and we hope to look at other kinds of types and relations that can be expressed even when the type theory lacks them, as quotient types can be by way of setoids. We also hope to tackle automation for proof repair directly in Cubical Agda, which is challenging because Cubical Agda lacks tools for building automation (except by reflection). We hope this will open the door to proof repair for more sophisticated classes of changes.
Data-Availabilty Statement
This paper comes with an artifact hosted on Zenodo containing the extension to Pumpkin Pi we wrote, as well as all the case studies described within the paper (Viola et al., 2025). Instructions for reproducing the results of the artifact can be found in Overview.md in the artifact. The source code for our extension to Pumpkin Pi and our case studies can be found on Github. Links to specific files in the Github source have been provided where relevant throughout the paper. These links are given as circled numbers, like 1.
Acknowledgements
We thank Amélia Liao, Reed Mullanix, Tom Jack, and the Cubical Agda Discord server for their help in using Cubical Agda. We thank also Carlo Angiuli for their early thoughts on this project. This material is based upon work supported by the Air Force Research Laboratory (AFRL), the Defense Advanced Research Projects Agency (DARPA), and the Naval Information Warfare Center Pacific (NIWC Pacific) under Contracts No. N66001-21-C-4023 and FA8750-24-C-B044. Any opinions, findings, and conclusions or recommendations expressed in this material are those of the authors and do not necessarily reflect the views of the AFRL, DARPA, and NIWC Pacific.